OpenStudy (anonymous):

multiply x^2 + 8x +15/x -4 * x^2 -16/ 2x +6

OpenStudy (anonymous):

factorize the numerators

OpenStudy (anonymous):

?

OpenStudy (anonymous):

factorize \[x^2 + 8x +15\]

OpenStudy (anonymous):

?how

sam (.sam.):

\[\frac{x^2 + 8x +15}{x -4} *\frac{ x^2 -16} {2x +6}\] \[\frac{(x+3)(x+5)}{x -4} *\frac{ (x-4)(x+4)} {2(x +3)}\] \[\frac{\cancel{(x+3)}(x+5)}{\cancel{x -4}} *\frac{\cancel{ (x-4)}(x+4)} {2\cancel{(x +3)}}\] \[\frac{(x+5)(x+4)}{2}\]